kingstonnnnn7678:

It took Jimena 89 minutes to run a 15-kilometer race last weekend. If you know that 1 kilometer equals 0.621 mile, how many minutes did it take Jimena to run 1 mile during the race? Round your answer to the nearest hundredth.

mw104969:

what are the answer choices

Convert:

15km*0.621m=9.315m 89/9.315=9.55448201825 minutes per mile Rounded=9.32 Answer: 9.32 minutes per mile

Convert:

I did that wrong i think ;-;

Convert:

9.55 minutes*

DK4L:

.15km*0.621m=9.315m 89/9.315=9.55 minutes per mile

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L .15km*0.621m=9.315m 89/9.315=9.55 minutes per mile \(\color{#0cbb34}{\text{End of Quote}}\) Tis what I said my dude :0

jhonyy9:

use the rule of 3 simple 89 min. to run 15 *0,621 miles x min. to run 1 mile x = 89/(15*0,621)= ?

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L .15km*0.621m=9.315m 89/9.315=9.55 minutes per mile \(\color{#0cbb34}{\text{End of Quote}}\) Tis what I said my dude :0 \(\color{#0cbb34}{\text{End of Quote}}\) o i didnt see that dude

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @DK4L .15km*0.621m=9.315m 89/9.315=9.55 minutes per mile \(\color{#0cbb34}{\text{End of Quote}}\) Tis what I said my dude :0 \(\color{#0cbb34}{\text{End of Quote}}\) o i didnt see that dude \(\color{#0cbb34}{\text{End of Quote}}\) No worries 0:

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert 9.55 minutes* \(\color{#0cbb34}{\text{End of Quote}}\) good job